OpenStudy (anonymous):

what fraction is equivilant to 5 3/4?

OpenStudy (anonymous):

5 3/4 is a fraction

OpenStudy (sosa4954):

17/4

OpenStudy (anonymous):

23/4 is the irregular fraction

OpenStudy (sosa4954):

no is it??

OpenStudy (anonymous):

17/4 = 4.25 = 4 1/4

OpenStudy (sosa4954):

oh

OpenStudy (anonymous):

23/4 = 5.75 = 5 3/4

OpenStudy (sosa4954):

oh ok lol

OpenStudy (anonymous):

wait you saw that?

OpenStudy (sosa4954):

saw what

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

ihere are the answer choices

sammixboo (sammixboo):

Well, what other fractions are equal to \(\rm \large \color{red}{\frac{3}{4}}\)?

OpenStudy (anonymous):

a. 15/4 b. 23/4 c. 8/4 d.19/4

OpenStudy (sosa4954):

burn i was right it is A

OpenStudy (anonymous):

who agrees?

OpenStudy (anonymous):

bye

OpenStudy (sosa4954):

I WAS RIGHT

OpenStudy (anonymous):

ok okay

sammixboo (sammixboo):

What is the improper fraction of \(\rm \large \color{red}{5\frac{3}{4}}\)? To do this you must multiply 4 by 5, so what is 5 x 4?

OpenStudy (anonymous):

20

sammixboo (sammixboo):

No you are not right. Please explain how you got your answer as well.

sammixboo (sammixboo):

Okay now you take the 20 and add 3, the numerator, so what is 20 + 3?

OpenStudy (anonymous):

23

sammixboo (sammixboo):

So now you take 23 and put it over the 4, our denominator, and that is your answer. \(\rm \large \color{red}{\frac{23}{4}}\)

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

i have another question

sammixboo (sammixboo):